OpenStudy (anonymous):

Find the slope of the line that passes through the points (0, 0) and (-2, -3). A) 2/3 B) 3/2 C) -3/2

OpenStudy (anonymous):

Find the slope first

OpenStudy (anonymous):

mhmmm

OpenStudy (mathstudent55):

The y-coordinates are: -3, and 0 Subtract -3 minus zero. What do you get?

OpenStudy (anonymous):

Which quadrant is it in? I, II, III, OR IV?

OpenStudy (anonymous):

From there you may continue finding the y cordinates

OpenStudy (mathstudent55):

The x-coordinates are -2 and 0. Subtract -2 minus 0. What do you get?

OpenStudy (anonymous):

Since the original was (0,0)

OpenStudy (anonymous):

Math student is correct in breaking down

OpenStudy (mathstudent55):

Then divide the difference of the y-coordinates by the difference of the x-coordinates to find the slope. Write it as a whole number or as a fraction.

OpenStudy (anonymous):

C? :)

OpenStudy (mathstudent55):

No. What did you get for -3 - 0? What did you get for -2 - 0?

OpenStudy (anonymous):

You have more steps to do after finding the slope before you can choose an answer, since you have to find an equation to go through one of the points.

OpenStudy (mathstudent55):

No. All the problem asks for is the slope of the line. You don't need to find the equation of the line.

OpenStudy (mathstudent55):

All you need to do is find the slope using the two given points. You can do this directly using the given information.

OpenStudy (anonymous):

2/3. I think? I've been trying to work this out on paper..

OpenStudy (anonymous):

otherwise, I am a complete idiot... :(

OpenStudy (mathstudent55):

You're not an idiot, but you need to pay attention. Once again, what is -3 - 0 ?

OpenStudy (anonymous):

(3/2) You flipped the fraction.

OpenStudy (anonymous):

@mathstudent55 @larryboxaplenty you both are very helpful, I just have a difficult time grasping concepts.. :|....

OpenStudy (mathstudent55):

-3 - 0 = -3 -2 - 0 = -2 The difference in y is -3. The difference in x is -2. Now divide the difference in y by the difference in x: (-3) / (-2) = 3/2 Answer is: slope is 3/2.

OpenStudy (anonymous):

You're so smart! @mathstudent55

OpenStudy (anonymous):

It helps me when people type out the steps like that!

OpenStudy (mathstudent55):

I went through it step by step asking you questions, but you never answered any of them. The key to learning a new concept is to break it down into simple steps. Each question I asked was a small step toward the solution.

OpenStudy (mathstudent55):

Right, that is the idea of taking one step at a time in finding the solution.

OpenStudy (mathstudent55):

You're welcome.